OpenStudy (anonymous):

What is the solution to the equation 3^(1-2x)=2? x≈-0.292 x≈-0.185 x≈0.185 x≈0.292 What is the solution to the equation 3^(2x)≈4? x=-0.631 x=-0.396 x=0.396 x=0.631

OpenStudy (maheshmeghwal9):

Take log!

OpenStudy (maheshmeghwal9):

both sides

OpenStudy (kropot72):

Take logs of both sides: \[(1-2x)\ln 3=\ln 2\] \[1-2x=\frac{\ln 2}{\ln 3}\] \[-2x=\frac{\ln 2}{\ln 3}-1\] -2x = 0.63 - 1 = -0.37 x = 0.185

OpenStudy (maheshmeghwal9):

Did u get?